Can someone please help me with these.
2. Solve by substitution or elimination method:
7x - 5y = 14
-4x + y = 27

3. Solve by substitution or elimination method:
-4x + 3y = 5
12x - 9y = -15

Answer by Alan3354(69443)

Can someone please help me with these.
2. Solve by substitution or elimination method:
7x - 5y = 14 eqn 1
-4x + y = 27 eqn 2
---------------------------
y = 4x+27 (from eqn 2)
Sub into eqn 1
7x - 5(4x+27) = 14
7x - 20x - 135 = 14
-13x = 149
x = -149/13
---------------
y = 4x+27
y = -596/13 + 27 = -596/13 + 351/13
y = -245/13
Sub to check:
7*(-149/13) - 5*(-245/13) ?? 14
-1043/13 + 1225/13 ?? 14
182/13 is 14, so it checks.
-----------------------------
3. Solve by substitution or elimination method:
-4x + 3y = 5
12x - 9y = -15
-----------------------
Divide the 2nd eqn by -3
-4x + 3y = 5
That's the same as eqn 1, so there's no unique solution. It's a straight line, or an infinite number of solutions.